The equation for molarity, M, is

M=n/V

where V is volume and n is the number of moles of solute.

A certain solution has a molarity of M = 7.23 mol/L and has a volume of V= 0.690 L . What is the value of n?
Answer #1

M = n ÷ V

7.23 mol / L = n ÷ 0.690 L

7.23 mol / L × 0.690 L = n

4.99 mol = n

So number of mole (n) = 4.99 mol